# Variable-Length Stop-Feedback Coding for Minimum Age of Incorrect Information

### v4

@@ Abstract.
conference: ; October 14–17, 2024; Athens, Greececcs: Networks Network performance analysisccs: Networks Network performance modelingccs: Mathematics of computing Coding theoryccs: Mathematics of computing Information theoryccs: Mathematics of computing Mathematical optimization
@@ 1. Introduction
Our work analyzes the minimum achievable average AoII in the non-asymptotic regime, where we investigate the impact of feedback time instances for variable-length stop-feedback (VLSF) codes. With VLSF coding, information is encoded into a theoretically infinite codeword, which is then segmented into packets. These packets are sequentially transmitted over the channel. Upon receiving a packet, the receiver attempts to decode the information, taking into consideration all previously received packets. Upon receiving a packet, the receiver attempts to decode it. Suppose the decoding is successful, meaning that the probability of error is less than or equal to a specified constant $\epsilon$. In that case, the receiver notifies the transmitter with an acknowledgement (ACK) feedback message. Otherwise, a negative acknowledgement (NACK) is sent to request more channel outputs.
@@
Several works have dealt with variable-length coding in the context of the AoI metric. Yet, most of them utilize the fixed-blocklength results of (Polyanskiy et al., 2010) to produce their results. The assumption that allows this utilization is that the decoder has perfect knowledge of decoding errors, even though the fundamental results of (Polyanskiy et al., 2010) do not adopt it. Examples of such works include (Sac et al., 2018) which analyzes variable-length codes, the paper (Yu et al., 2020) that approximates the optimal blocklength for vehicular networks, the paper (Xie et al., 2021) for multicast networks under energy constraints, and (Devassy et al., 2018) that demonstrates the effect of blocklength on the violation probabilities of delay and peak AoI. Notably, the latter is extended for proper variable-length codes in (Devassy et al., 2019), by utilizing the results of (Polyanskiy et al., 2011).
To the best of our knowledge, this is the first work that analyzes the AoII with VLSF codes and the first work in the AoI/AoII literature that considers the feedback sequence as part of the optimization problem.
@@ 2.2. Feedback Sequence
In practical communication systems, the maximum number of packets or channel uses is constrained (Ahmed et al., 2021). We achieve this by requiring the following equality,
@@ 2.3. Source Model
(8) $\mathbf{P}_{ii}^{t}>\mathbf{P}_{ij}^{t},\quad\forall\ i\neq j,t\in\mathbbm{N},$
@@ 3. Mathematical Formulation
Thus, in the absence of successful decodings, the AoII progresses as a Markov chain, which is illustrated in Fig. 3.
@@
Next, let the event $D_{t+1}$ occur. Let $b_{t+1}$ denote the number of symbols sent for the sample being transmitted at time $t+1$, and also let $l_{t+1}$ denote the length of the current packet. Note that the transmitter sends only the freshest source values, discarding the old sample and starting a new transmission if needed, and thus the decoded value equals $\hat{X}_{t+1}=X_{t+1-l_{t+1}}$. In this context, the decoded value is correct if the source value has not changed, i.e. $X_{t+1}=X_{t+1-l_{t+1}}$, which happens with probability $\mathbf{P}_{ii}^{l_{t+1}}$. Hence,
(16) $P(\delta_{t+1}=0\mid D_{t+1})=\mathbf{P}_{ii}^{l_{t+1}},$
@@
The MDP is similar to the AoII-optimal MDP (Def. 1), with the exception that $\mathbf{P}_{ii}^{l}\doteq 1\ \forall\ l\in\mathbbm{N}$ and $d_{\text{feedback}}\doteq 0$.
@@
An exact solution of the Bellman equations is generally hard to find but can be approximated with the relative value iteration (RVI) algorithm (Krishnamurthy, 2016). Particularly, the RVI approximates the value function iteratively and, when it converges, it finds the true value function. In particular, let $V_{\tau}(S)$ denote the estimation of $V(S)$ at iteration $\tau\in\mathbbm{N}$. Next, define the Bellman operator,
@@
(22) $V_{\tau+1}(S)=TV_{\tau}(S)-TV_{\tau}(S_{0}),\quad\forall\ S\in\mathcal{S},\ \tau\in\mathbbm{N}\,.$
@@ 5. Conclusions
Acknowledgements
The work was partially funded by the project POLAR (FORTH Synergy Grant 2021) and the project ARGOS-CDS (European Commission, Grant Agreement number: 101094354), and partially by the Engineering and Physical Sciences Research Council (EP/L016656/1) and the University of Bristol.

### v5

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The Twenty-fifth International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing; October 14–17, 2024; Athens, Greecebooktitle: The Twenty-fifth International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing (MOBIHOC ’24), October 14–17, 2024, Athens, Greecedoi: 10.1145/3641512.3690165isbn: 979-8-4007-0521-2/24/10conference: ; October 14–17, 2024; Athens, Greececcs: Networks Network performance analysisccs: Networks Network performance modelingccs: Mathematics of computing Coding theoryccs: Mathematics of computing Information theoryccs: Mathematics of computing Mathematical optimizationfootnotetext: This is the extended version of the paper presented at the MobiHoc ’24 Workshops.
Published version: https://doi.org/10.1145/3641512.3690165
@@ 1. Introduction
Our work analyzes the minimum achievable average AoII in the non-asymptotic regime, where we investigate the impact of feedback time instances for variable-length stop-feedback (VLSF) codes. With VLSF coding, information is encoded into a theoretically infinite codeword, which is then segmented into packets. These packets are sequentially transmitted over the channel. Upon receiving a packet, the receiver attempts to decode the information, taking into consideration all previously received packets. Suppose the decoding is successful, meaning that the probability of error is less than or equal to a specified constant $\epsilon$. In that case, the receiver notifies the transmitter with an acknowledgement (ACK) feedback message. Otherwise, a negative acknowledgement (NACK) is sent to request more channel outputs. Noticeably, this communication setting is very similar to hybrid automatic repeat request with incremental redundancy (HARQ-IR).
@@
Several works have dealt with variable-length coding with finite blocklength in the context of the AoI metric. Yet, most of them utilize the fixed-blocklength results of (Polyanskiy et al., 2010) in a simple ARQ setting assuming the decoder has perfect knowledge of decoding errors. Examples of such works include (Sac et al., 2018) which analyzes variable-length codes, the paper (Yu et al., 2020) that approximates the optimal blocklength for vehicular networks, the paper (Xie et al., 2021) for multicast networks under energy constraints, and (Devassy et al., 2018) that demonstrates the effect of blocklength on the violation probabilities of delay and peak AoI. Notably, the latter is extended for VLSF coding in (Devassy et al., 2019). All of the above works assume instantaneous feedback. Moreover, they focus on equi-length packets, with the exception of (Devassy et al., 2019) which also considers the average blocklength in VLSF coding.
Very few papers consider the feedback sequence in retransmission schemes. We note (Arafa et al., 2019), which draws the AoI-optimal feedback sequence for stationary policies with instantaneous feedback. In (Munari and Badia, 2022), the presence of feedback incurs a reduction in the number of transmissions, leading to insightful results. Our work differentiates itself by directly considering the feedback delay.
To the best of our knowledge, this is the first work to analyze the AoII with VLSF coding. It is also the first to optimize the AoII with positive feedback delay when monitoring sources with small cardinality. Additionally, this is the first time in the AoII research that the feedback sequence is included as part of the optimization problem.
@@ 2.2. Feedback Sequence
In practical communication systems, the maximum number of packets or channel uses is constrained (et al., 2021). We achieve this by requiring the following equality,
@@ 2.3. Source Model
(8) $\mathbf{P}_{ii}^{t}>\mathbf{P}_{ij}^{t},\quad\forall\ i\neq j,t\in\mathbb{N},$
@@ 3. Mathematical Formulation
Thus, in the absence of successful decodings, the AoII progresses as a Markov chain, illustrated in Fig. 3.
@@
Next, let the event $D_{t+1}$ occur. Let $b_{t+1}$ denote the number of symbols sent for the sample being transmitted at time $t+1$, and also let $l_{t+1}$ denote the corresponding packet length. Note that the transmitter sends only the freshest source values, discarding the old sample and starting a new transmission if needed, and thus the decoded value equals $\hat{X}_{t+1}=X_{t+1-l_{t+1}}$. In this context, the decoded value is correct if the source value has not changed, i.e. $X_{t+1}=X_{t+1-l_{t+1}}$, which happens with probability $\mathbf{P}_{ii}^{l_{t+1}}$. Hence,
(16) $P(\delta_{t+1}=0\mid D_{t+1})=P(X_{t+1}=X_{t})=\mathbf{P}_{ii}^{l_{t+1}},$
@@
The MDP is similar to the AoII-optimal MDP (Def. 1), with the exception that $\mathbf{P}_{ii}^{l}\doteq 1\ \forall\ l\in\mathbb{N}$ and $d_{\text{feedback}}\doteq 0$.
@@
An exact solution of the Bellman equations is generally hard to find but can be approximated with the relative value iteration (RVI) algorithm (Krishnamurthy, 2016). Particularly, the RVI approximates the value function iteratively and, when it converges, it finds the true value function. In particular, let $V_{\tau}(S)$ denote the estimation of $V(S)$ at iteration $\tau\in\mathbb{N}$. Next, define the Bellman operator,
@@
(22) $V_{\tau+1}(S)=TV_{\tau}(S)-TV_{\tau}(S_{0}),\quad\forall\ S\in\mathcal{S},\ \tau\in\mathbb{N}\,.$
@@ 5. Conclusions
Acknowledgements.
This work was partially funded by the project POLAR (FORTH Synergy Grant 2021) and by the European Union – NextGenerationEU under the framework of the National Recovery and Resilience Plan Greece 2.0, grant (Project: Smart Cities, TAEDR-0536642), and partially by the Engineering and Physical Sciences Research Council (EP/L016656/1) and the University of Bristol.